OpenStudy (anonymous):

Your piggy bank has a total of 20 coins in it; some are dimes and some are nickels. If you have a total of $1.25, how many nickels do you have? 15 10 5 20

OpenStudy (hihi67):

Okay, let's start with ten nickels.

OpenStudy (hihi67):

That would be a total of 50 cents.

OpenStudy (hihi67):

You still have ten coins left, and they have to be dimes.

OpenStudy (anonymous):

how did you do that

OpenStudy (anonymous):

i meant 15 coins not 20

OpenStudy (hihi67):

And ten times ten is 100 cents, so a dollar. And a dollar plus 50 cents is 1.50, which is too much.

OpenStudy (hihi67):

???

OpenStudy (anonymous):

my problem on my homeowrk says this a cash register contains 15 coins in dime and nickels. the total value is 1.25. how many coins of each typer are there

OpenStudy (hihi67):

Oh, okay.

OpenStudy (hihi67):

Then I'll start over.

OpenStudy (hihi67):

Just pick a random number of nickels to start with...let's say 7.

OpenStudy (hihi67):

If you have 7 nickels, how much money is that?

OpenStudy (anonymous):

35 cents

OpenStudy (hihi67):

Good, and how many coins do we have left that have to be dimes?

OpenStudy (anonymous):

8 coins left...

OpenStudy (hihi67):

Good, and since they have to be dimes, how much money is that?

OpenStudy (anonymous):

80 cents?

OpenStudy (hihi67):

Sorry, my cat just walked on the keyboard and deleted a bunch of stuff. And yes, your answer is right.

OpenStudy (hihi67):

And 80 + 35 is?

OpenStudy (anonymous):

45

OpenStudy (hihi67):

Uh, no. It's 1.15

OpenStudy (hihi67):

So, not enough.

OpenStudy (hihi67):

Let's try again with 8 nickels.

OpenStudy (hihi67):

8 times 5 is 40. 7 times 10 is 70. 40+70 is 1.30

OpenStudy (anonymous):

still too much

OpenStudy (hihi67):

Yeah, what I don't understand, is with 7 nickels it's too little, and with 8 it's not enough...

OpenStudy (anonymous):

ya. thanks anyway. i get it somehow. it suppose be into a equation

OpenStudy (hihi67):

Oh, sorry, should I help you with an equation?

OpenStudy (texaschic101):

d + n = 20 -->d = 20 - n .10d + .05n = 1.25 now sub 20 - n in for d in the second equation and solve for n

OpenStudy (hihi67):

Like that made sense to her.

OpenStudy (hihi67):

d is dimes n is nickels

OpenStudy (hihi67):

a dime is a tenth of a dollar so therefore .10

OpenStudy (hihi67):

a nickel is a twentieth so therefore .05

OpenStudy (hihi67):

d and n are variables, so they can be any number. if you multiply d with .10 it means for example 5 dimes times 0.10, which is .50, or 50 cents.

OpenStudy (hihi67):

Do you get it @bs?

OpenStudy (hihi67):

ugh, she's offline

OpenStudy (texaschic101):

don't you hate that....do all that explaining and she is gone

OpenStudy (hihi67):

Yeah, ikr

OpenStudy (hihi67):

lol, but now we both got medals! :)

OpenStudy (texaschic101):

yes we do :)

OpenStudy (austinl):

@texaschic101 's way is MUCH easier. A system of equations. \(\Large{d+n=20\\ .10d + .05n = 1.25}\) This is how I would set up the problem, and I am fairly confident that this is how the problem was intended to be solved.

OpenStudy (texaschic101):

that is how I would set it up also :)